Skip to end of metadata
Go to start of metadata

AHATS status 8/11/08

Staff: Militzer, Oncley, Nguyen, Verstraete with ISS.

Temps: 97F/57F yesterday

Activity highlights:

  • now in array4.
  • See profile note, some outages last night, board problem
  • Logbook Wiki and NCAR Web server is still down.
    [The work on a sql server update on saturday caused the problem.
    The vendor/ncar staff expect a patch by midnight tonight will allow
    'logbook' to be back up by tomorrow morning.]

https://wiki.ucar.edu/display/ahatslogbook
...but wiki.ucar.edu has been down this weekend...

Good wind direction:
array#4 (aug 9 18:00 - present) has 11.42 hours (.42 unstable, 11 stable).

Local data storage: (swapped pressure1/2 sticks,upwind pocketec yesterday)
------------------
upwind:/dev/sda1 57685532 1594044 56091488 3% /var/tmp/usbdisk
downwind1:/dev/sda1 57685532 9538676 48146856 17% /var/tmp/usbdisk
downwind2:/dev/sda1 57685532 9744300 47941232 17% /var/tmp/usbdisk
profile:/dev/sda1 57685532 7100640 50584892 12% /var/tmp/usbdisk
pressure1:/dev/sda1 3940812 398016 3542796 10% /var/tmp/usbdisk
pressure2:/dev/sda1 3940812 401564 3539248 10% /var/tmp/usbdisk

aster:/dev/sdb1 721075720 460821376 260254344 64% /media/isff2
aster:/dev/sdb1 1922890480 640618164 1184595112 36% /media/isff15

Pressure:
--------
(+/- = ~1 std deviation among variables at the same height)

P.bedard had a pinched hose overnight so data is bogus until fixed today @~10pdt.

'p.ref' is measuring AHATS ref p with the 202BG connected between the
AHATS and CHATS references.
P.bedard now is the portable Paroscientific standard with reasonable resolution, but
low sampling rate connected to the Bedard pressure port now installed on boom 13b.

p: ok, +/- 0.03 mb plot scale too large to tell
p'p': ok, +/- 0.0005 mb^2 small but perhaps okay -- only stable data to look at
w'p': ok, +/- 0.002 m/s mb
t: ok, 1 degC
Pref: ok, reconnected to volume at start of this config.
P: ok

Profile: ,select.p
--------
Diamond board problems overnight and today. Lost ~8hours.
Need to swap new 8P board with an old version 8M, but none available.
Started crontab to 'adn / aup' every hour to force a restart. Seems to be helping
and we've gotten data since doing this ~11pdt.
Heights of sensors are the same and array3.

T: ok
RH: sensor that is at 5.8m is up to 2% off. Didn't change throughout array #3 period,
now is not important in array #4.

h2o: normal, 2 g/m^3 offset from dat("Q"), except in afternoon.
w'h2o': ok, 0.025 m/s g/m^3 at midday
co2: ok, 14-17 mmol/m^3 (some nights have large values, some don't)
w'co2': ok, -0.01 m/s mmo/m^3 at midday

diag: ok? spikes in 4.8m.
samples.sonic: ok
spd: ok?, Swapped sonic at 8m.
Replaced 720 with 733. 720 had been reporting a bit too high. Had been at
7m, and when we swapped it to 8m the high reading showed up.
dir: ok
w: ok
tc: ok, still strange 7m/8m ordering.

-----------------------
-----------------------
Note: From here on the specific sensor comments are removed because we
began a rerun of all of the covars which caused display/naming issues with
running the 'daily.qq' plots. From the brief looks, everything was up and
running in down1,2, pressure1,2 upwind and profile.